1 train leaves a location at 12:00pm travelling
At 40 mph. The second train leaves from the same location
2 hours later travelling at 60mph. If both trains are travelling on parallel
Tracks, what time will the 2nd train pass the 1st train?

Answer by Maths68(1474)

Car A Data
==========
Speed = 40mph
Time = t hour
Distance covered = x miles
Distance = speed*time
x=40*t
x=40t.............(1)
Car B Data
==========
Speed = 60mph
Time =t-2(hour)
Distance covered = x
Distance = Speed * time
x=60*(t-2)
Both cars will cover same distance
40t=60*(t-2)
40t=60t-120
40t-60t=-120
-20t=-120
-20t/-20=-120/-20
t=6 hours
Car A travel time 6 hours and Car B travel time 4 hours to cover the same distance.
After 4 hours of leaving same place Car B will catch Car A
Check
========
Car A Data
==========
Speed = 40mph
Time = 6 hour
Distance covered = x miles
Distance = speed*time
x=40*6
x=240 miles
Car B Data
==========
Speed = 60mph
Time =t-2=6-2= 4(hour)
Distance covered = x
Distance = Speed * time
x=60*4
x=240 miles